OpenStudy (anonymous):

Use Euler's method with step size 0.1 to estimate y(1.5), where y(x) is the solution of the initial-value problem y' = 3y + 2xy, y(1) = 1.

OpenStudy (anonymous):

Please someone help, I will give 10 medals .

OpenStudy (anonymous):

This problem is due in 10 mintues, please someone help!

OpenStudy (anonymous):

what class is this?

OpenStudy (anonymous):

Calc 2

OpenStudy (anonymous):

you need to construct a table first

OpenStudy (anonymous):

Start filling out a table:\[\begin{array}{c|c|c|c|c} n&x_n&y_n&y'(x_n,y_n)\\ \hline 0&\color{green}1&\color{red}1&3(\color{red}1)+2(\color{green}1)(\color{red}1)=\color{blue}5\\ \hline 1&\color{green}1+0.1=1.1&\color{red}1+\color{blue}5(0.1)=1.5\\ \hline 2&\\ \hline 3&\\ \hline 4&\\ \hline 5& \end{array}\]

OpenStudy (anonymous):

The next step would be \[\begin{array}{c|c|c|c|c} n&x_n&y_n&y'(x_n,y_n)\\ \hline 0&1&1&5\\ \hline 1&\color{green}{1.1}&\color{red}{1.5}&2(\color{red}{1.5})+3(\color{green}{1.1})(\color{red}{1.5})=\color{blue}{7.95}\\ \hline 2&\color{green}{1.1}+0.1=1.2&\color{red}{1.5}+\color{blue}{7.95}(0.1)=2.295\\ \hline 3&\\ \hline 4&\\ \hline 5&&&- \end{array}\] and so on. Follow the color-coding if you like, but it's better to get an idea of WHY we're doing the calculations we're doing.